# Reciprocal Polynomials with Zeros on the Unit Circle and Derivatives of Chebyshev Polynomials of the Second Kind

### v1

@@ 1 Introduction
  • let $\left\{\nu_{j}\right\}_{j=1}^{\left[\frac{N-3}{2}\right]}$ be the set of positive roots of the equation $U^{\prime}_{N-2}(x)=0$ and let $\gamma_{j}=1-2(\nu_{j})^{2}$; also, let $\left\{\mu_{j}\right\}_{j=1}^{\left[\frac{N-3}{2}\right]}$ be the set of positive roots of the equation $U_{N-2}(x)=0$ and let $\beta_{j}=1-2(\mu_{j})^{2}$ when $N$ is odd; then the following factorization formulas are valid for $q(z)$ (for $p(z)$ there are similar formulas):
@@
      $\kappa=-\frac{N}{N-2}$, $N$ even
@@
      $\kappa=\frac{N}{N-2}$, $N$ even
@@ 4 Extremal properties of polynomials with zeros on the unit circle
If all zeros of a polynomial $P(z)$ are located on the unit circle, then all of the zeros of the polynomial $P^{\prime}(z)$ lie inside the disc $\overline{\mathbb{D}}$. Since the zeros of the derivative of a polynomial lie inside the closed convex hull of the zeros of this polynomial, then all the zeros of the polynomial $P^{(s+1)}(z)$ must lie inside the disc $\overline{\mathbb{D}}$. From Vieta’s theorem it follows that the coefficients of the polynomial $\sum_{j=0}^{s}(-1)^{j}\gamma_{j}\frac{N!}{(N+s+1)!}\frac{(N+s+1-j)!}{(N-j)!}z^{s-j}$ must satisfy the constraints: $|\gamma_{j}|\frac{N!}{(N+s+1)!}\frac{(N+s+1-j)!}{(N-j)!}\leq{s\choose s-j}$ for $j=1,\ldots,s$, which are equivalent to conditions (13). Generally speaking, it is impossible to reduce the validity of even one of the inequalities (13), as shown by the example from the Corollary 3.3. The theorem is proved. ∎
@@
Moreover, without additional assumptions on the coefficients, neither the sides of the parallelepiped $\hat{\Pi}$ nor the edges of the cube $\tilde{\Pi}$ can be reduced or increased. For the latter case, an example would be a polynomial $P(z)=\left(1-z^{N+3}\right)+\gamma_{2}\left(z^{2}-z^{N+1}\right)$. Then, $\gamma_{2}\leq 1$ for odd $N$ and $\gamma_{2}\geq-1$ for even $N$ [5].
@@ 5 Some examples
However, this criterion can only be used to check whether all zeros of the polynomial (12) belong to the unit circle when all coefficients of the polynomial are exactly known. To construct sets in the space of coefficients for which all zeros of the polynomial (12) belong to the unit circle, condition A must be supplemented with other conditions, which can be obtained, for example, from Theorem 1.1.
@@
    for odd N1N-1, and $U_{\pi}=\left\{\gamma_{1}\ :\ \gamma_{1}=-1\right\}$ for even N1N-1. Thus, a set in the coefficient space of the polynomials (12) for which all zeros of the polynomial belong to the unit circle is defined by many $\left\{\gamma_{1}\ :\ \left|\gamma_{1}\right|\leq\frac{N+2}{N}\right\}$ for even $N$ and many $\left\{\gamma_{1}\ :\ -1\leq\gamma_{1}\leq\frac{N+2}{N}\right\}$ for odd $N$.

### v2

@@ 1 Introduction
  • let $\left\{\nu_{j}\right\}_{j=1}^{\left[\frac{N-3}{2}\right]}$ be the set of positive roots of the equation $U^{\prime}_{N-2}(x)=0$ and let $\gamma_{j}=1-2(\nu_{j})^{2}$; also, let $\left\{\mu_{j}\right\}_{j=1}^{\left[\frac{N-3}{2}\right]}$ be the set of positive roots of the equation $U_{N-2}(x)=0$ and let $\beta_{j}=1-2(\mu_{j})^{2}$ when NN is odd; then the following factorization formulas are valid for $q(z)$ (for $p(z)$ there are similar formulas):
@@
      $\kappa=-\frac{N}{N-2}$, NN even
@@
      $\kappa=\frac{N}{N-2}$, NN even
@@ 4 Extremal properties of polynomials with zeros on the unit circle
If all zeros of a polynomial $P(z)$ are located on the unit circle, then all of the zeros of the polynomial $P^{\prime}(z)$ lie inside the disc $\overline{\mathbb{D}}$. Since the zeros of the derivative of a polynomial lie inside the closed convex hull of the zeros of this polynomial, then all the zeros of the polynomial $P^{(s+1)}(z)$ must lie inside the disc $\overline{\mathbb{D}}$. From Vieta’s theorem it follows that the coefficients of the polynomial $\sum_{j=0}^{s}(-1)^{j}\gamma_{j}\frac{N!}{(N+s+1)!}\frac{(N+s+1-j)!}{(N-j)!}z^{s-j}$ must satisfy the constraints: $|\gamma_{j}|\frac{N!}{(N+s+1)!}\frac{(N+s+1-j)!}{(N-j)!}\leq{s\choose s-j}$ for $j=1,\ldots,s$, which are equivalent to conditions (13). Generally speaking, it is impossible to reduce the validity of even one of the inequalities (13), as shown by the example from the Corollary $3.$ The theorem is proved. ∎
@@
Moreover, without additional assumptions on the coefficients, neither the sides of the parallelepiped $\hat{\Pi}$ nor the edges of the cube $\tilde{\Pi}$ can be reduced or increased. For the latter case, an example would be a polynomial $P(z)=\left(1-z^{N+3}\right)+\gamma_{2}\left(z^{2}-z^{N+1}\right)$. Then, $\gamma_{2}\leq 1$ for odd NN and $\gamma_{2}\geq-1$ for even NN [5].
@@ 5 Some examples
However, this criterion can only be used to check whether all zeros of the polynomial (12) belong to the unit circle when all coefficients of the polynomial are exactly known. To construct sets in the space of coefficients for which all zeros of the polynomial (12) belong to the unit circle, condition A must be supplemented with other conditions, which can be obtained, for example, from Theorem $1.$
@@
    for odd N1N-1, and $U_{\pi}=\left\{\gamma_{1}\ :\ \gamma_{1}=-1\right\}$ for even N1N-1. Thus, a set in the coefficient space of the polynomials (12) for which all zeros of the polynomial belong to the unit circle is defined by many $\left\{\gamma_{1}\ :\ \left|\gamma_{1}\right|\leq\frac{N+2}{N}\right\}$ for even NN and many $\left\{\gamma_{1}\ :\ -1\leq\gamma_{1}\leq\frac{N+2}{N}\right\}$ for odd NN.